Sarah spent 2 fifths of her paycheck on rent. If her rent is $80 how much was she paid?
Answer by Theo(13342)

x equals her rent.
2/5*x equals the rent.
2/5*x = 80
solve for x.
multiply both sides by (5/2) to getr:
x = 80 * (5/2) which becomes:
x = 400/2 which becomes:
x = 200
she was paid 200
2/5 * 200 = 400/5 = 80